OpenStudy (anonymous):

How do you solve one equation with three variables? −2x+3y+2z=6

OpenStudy (anonymous):

You can only solve for a single variable as an expression of the other 2 variables

OpenStudy (anonymous):

Use cramer's rule: http://www.purplemath.com/modules/cramers.htm

OpenStudy (anonymous):

Diofeher that is only if you are given 3 different equations and what they equal, here there is only 1 equation with 3 unknowns

OpenStudy (anonymous):

manumben, how would I go about doing that? I'm learning system of equations, so I don't know exactl what to do. iknow i have to find it in terms of s1, s2, and s3 i believe

OpenStudy (anonymous):

With the given information you can't. you need more to the problem

OpenStudy (anonymous):

I'm not sure what you're asking but this is one possibility:\[−2x+3y+2z=6~~~\implies~~~-2x=6-3y-2z~~~\implies~~~\color{red}{x=-3+{3\over2}y+z}\]\[−2x+3y+2z=6~~~\implies~~~~~~~3y=6+2x-2z~~~\implies~~~\color{red}{y=2+{2\over3}x-{2\over3}z}\]\[−2x+3y+2z=6~~~\implies~~~~~~~2z=6+2x-3y~~~\implies~~~\color{red}{z=3+x-{3\over2}y}\]

OpenStudy (anonymous):

@paki123 , Hope this helps!